# A Localized Orthogonal Decomposition method for heterogeneous mixed-dimensional problems

### v1

@@ 1. Introduction
From a numerical perspective, solving mixed-dimensional PDEs presents several challenges. Classical finite element methods require meshes that resolve the complex geometry of embedded features, which can become intractable for a large number of interfaces. The situation is further complicated by the presence of strongly heterogeneous and oscillatory coefficients, which must also be resolved by the considered meshes. This situation is typical in porous media flow where the permeability varies rapidly in the bulk regions between interfaces. This paper focuses precisely on such a setting: we study a mixed-dimensional, elliptic diffusion problem with highly heterogeneous coefficients.
Many numerical methods have been developed for solving partial differential equations on surfaces. A foundational contribution to the finite element approximation of the Laplace–Beltrami equation was given in [13], where the method is formulated on a polyhedral approximation of the surface (see also the comprehensive overview in [10]). Alternative approaches include trace-based methods [5, 3, 33], in which surface functions are represented as traces of functions defined in a higher-dimensional ambient space. Finite element methods for coupled bulk–interface problems have also been the subject of extensive research; see, for example, [17, 29]. More recently, a fitted finite element method has been introduced in [23], with particular emphasis on establishing Poincaré-type inequalities for the coupled bulk–interface model.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce the mixed-dimensional model problem. A prototypical multiscale method that achieves optimal convergence rates under minimal structural assumptions is presented in Section 3. In Section 4, we establish the exponential decay of the corresponding prototypical basis functions, motivating their localization, which is discussed in Section 5. Based on the resulting localized basis functions, a practical multiscale method is developed in Section 6. In Section 7 we extend the methodology to allow for interfaces that vary of the fine computational mesh. Finally, in Section 8, we present a series of numerical experiments that support our theoretical findings.
@@ 2.1. Geometry
To formulate assumptions on the topology of the problem, we define, for each codimension $c$, a topological subspace $X^{c}\subset\mathbb{R}^{d}$. For codimension $c=0$, we set $X^{0}=\Omega$, and for codimensions $c>0$, we define $X^{c}=X^{c-1}\setminus\Omega^{c-1}.$ Standard topological concepts such as openness, closedness, and density are understood with respect to the subspace topology on $X^{c}$. The closure, boundary, and interior operators are denoted by $\overline{\cdot}$, $\partial\cdot$, and $\operatorname{int}(\cdot)$, respectively. We assume that each subdomain segment $\Omega^{c}_{\ell}$, for $\ell\in L_{c}$, is open in $X^{c}$, and that their union $\Omega^{c}$ is dense in $X^{c}$. Furthermore, we suppose that for any pair $(i,j)\in L_{c}\times L_{c+1}$, we either have the inclusion $\Omega^{c+1}_{j}\subseteq\partial\Omega^{c}_{i}$ or it holds $\Omega^{c+1}_{j}\cap\overline{\Omega^{c}_{i}}=\emptyset$. This condition allows us to define an adjacency relation $E^{c}$ between segments of codimensions $c$ and $c+1$ as
@@ 2.2. Function spaces
where the space $V^{0}$, defined on the bulk domain of codimension 0, is broken along the interfaces between subdomain segments. Specifically,
@@ 2.3. Problem formulation
$-\nabla^{1}\cdot\left(A^{1}\nabla^{1}u^{1}\right)-\sum_{(i,j)\in E_{0}}B^{1}\!\left(u^{0}_{i}-u^{1}_{j}\right)$ $=f^{1},$ $\text{in }\Omega^{1},$
@@
The coercivity proof employs an iterative procedure, starting by bounding norms on subdomain segments far from the domain boundary and progressively extending these bounds to neighboring segments until reaching those where a Friedrichs-type inequality applies; see [23, Lem. 1] for details. The boundedness can be proved using a standard trace inequality; see [23, Lem. 2]. ∎
@@ 3. Prototypical method
This section presents a prototypical multiscale method that achieves optimal-order approximations without any pre-asymptotic effects, under minimal structural assumptions on the coefficients. The LOD method is typically defined with respect to a coarse mesh of the domain. In the present mixed-dimensional setting, coarse meshes are required for both the bulk domain and the interface domain, denoted by $\mathcal{T}_{H}^{0}$ and $\mathcal{T}_{H}^{1}$, respectively. They are henceforth referred to as the bulk mesh and the interface mesh, respectively. Note that the considered coarse meshes are not required to resolve the fine-scale variations of the coefficients. In general, there is considerable flexibility in how these meshes are chosen. Even complex-shaped elements representing fractures can be accommodated, provided the elements are non-degenerate and a local element-wise Poincaré inequality holds.
@@
(3.5b) $\qquad\qquad b(\tilde{\varphi}_{T^{c}},\mu)$ $=\quad$ $\mu_{T^{0}},$ $\text{ for all }\mu\in\mathbb{R}^{N},$
@@
(3.7) $\begin{split}\|\nabla^{c}v_{T^{c}}\|_{L^{2}(T^{c})}&\lesssim H^{-1}\|v_{T^{c}}\|_{L^{2}(T^{)}}\lesssim H^{(d-c)/2-1},\quad c\in\{0,1\}.\end{split}$
@@
We begin by establishing the well-posedness of problem 3.5 by verifying the inf–sup stability of the bilinear form $b$. Specifically, we show that there exists a constant $C>0$ such that
@@
Given any $\mu\in\mathbb{R}^{N}$, we construct a function $v\in V_{0}$ component-wise as
@@
Having introduced the prototypical basis functions, we can define a projection operator $\mathcal{R}\colon V\to\tilde{V}_{H}$ which preserves the QOI from 3.1 by
@@
This operator coincides with the $a$-orthogonal projection onto $\tilde{V}_{H}$, since for any $v\in V_{0}$ and $w\in\tilde{V}_{H}$ it holds that $a(v-\mathcal{R}v,w)=0$, noting that $v-\mathcal{R}v\in W$. This orthogonality relation also implies the continuity of $\mathcal{R}$, i.e.,
@@ 4. Exponential decay of prototypical basis
where $\mathcal{T}^{c}(z^{c})$ denotes the set of elements in $\mathcal{T}_{H}^{c}$ sharing the node $z^{c}$, and $\Lambda_{z^{c}}$ is the classical hat function associated with node $z^{c}$. For boundary nodes, the corresponding nodal values of $\mathcal{I}_{H}^{c}v$ are set to zero to conform to the prescribed Dirichlet boundary conditions. This operator satisfies for $c\in\{0,1\}$ the following local approximation and stability estimates:
(4.2) $\|v-\mathcal{I}_{H}^{c}v\|_{L^{2}(T^{c})}+\|\nabla^{c}\mathcal{I}_{H}^{c}v\|_{L^{2}(T^{c})}\lesssim\|\nabla^{c}v\|_{L^{2}(\mathsf{N}^{c}(T^{c}))},\quad v\in V^{c},\;T^{c}\in\mathcal{T}_{H}^{c}$
@@
where, $a_{T^{0}}$ and $b_{T^{0}}$ denote, for any $T^{0}\in\mathcal{T}_{H}^{0}$, suitable restrictions of the bilinear forms $a$ and $b$, satisfying the properties
@@
(4.7) $\begin{split}a_{T^{0}}(v,w)&=({A^{0}\nabla^{0}v^{0}},{\nabla^{0}w^{0}})_{T^{0}}+\sum_{T^{1}\subset\partial T^{0}}\frac{1}{n(T^{1})}({A^{1}\nabla^{1}v^{1}},{\nabla^{1}w^{1}})_{T^{1}}\\ &\qquad+\sum_{T^{1}\subset\partial T^{0}}({B^{1}(v^{0}|_{T^{0}}-v^{1})},{w^{0}|_{T^{0}}-w^{1}})_{T^{1}},\end{split}$
@@
(4.9) $\begin{split}\|\varphi\|_{V(S)}^{2}&\coloneqq\|\nabla^{0}\varphi^{0}\|_{L^{2}(S^{0})}^{2}+\|\nabla^{1}\varphi^{1}\|_{L^{2}(S^{1})}^{2}+\hskip-5.38193pt\sum_{\begin{subarray}{c}\{T^{1}\subset\partial T^{0}\,:\,\\ T^{0}\subset S^{0},\,T^{1}\subset S^{1}\}\end{subarray}}\hskip-5.38193pt\|v^{0}|_{T^{0}}-v^{1}\|_{L^{2}(T^{1})}^{2}.\end{split}$
@@
(4.10) $\|\mathcal{K}_{T^{0}}v\|_{V(\Omega\setminus\mathsf{N}^{\ell}(T^{0}))}\lesssim\exp(-c\ell)\|\mathcal{K}_{T^{0}}v\|_{V(\Omega)},\quad v\in V_{0}.$
@@
For a given $v\in V_{0}$, define $\varphi\coloneqq\mathcal{K}_{T}v$. Introduce a cut-off function $\eta\in V_{H}^{0}$, defined on the bulk domain, and uniquely characterized by the following properties:
@@
$|\lambda_{K^{0}}|$ $=|b(\lambda,b_{K^{0}})|=|a(\varphi,b_{K^{0}})|=|(A^{0}\nabla^{0}\varphi^{0},\nabla^{0}v_{K^{0}})_{L^{2}(\Omega^{0})}$
@@
(4.18) $\sum_{K^{1}\subset R_{T^{0}}^{1}}H^{-d+3}|\lambda_{K^{1}}|^{2}\lesssim\|\nabla^{0}\varphi^{0}\|^{2}_{L^{2}(\mathsf{N}^{0}(R_{T^{0}}^{0}))}+\|\nabla^{1}\varphi^{1}\|^{2}_{L^{2}(R_{T^{0}}^{1})}.$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}\leq C\|\varphi\|^{2}_{V(\mathsf{N}^{\ell+1}(T^{0})\setminus\mathsf{N}_{\ell-2}(T^{0}))}.$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}$ $\leq\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell-1}(T^{0}))}\leq C\|\varphi\|^{2}_{V(\mathsf{N}_{\ell}(T^{0}))\setminus\mathsf{N}_{\ell-3}(T^{0}))}$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}^{\ell}(T^{0}))}\leq\gamma\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}^{\ell-3}(T^{0}))}\leq\gamma^{\lfloor\ell/3\rfloor}\|\varphi\|^{2}_{V(\Omega)},$
@@ 5. Localization
$\|\mathcal{K}^{\ell}_{T^{0}}v\|_{V(\Omega\setminus\mathsf{N}^{m}(T^{0}))}\lesssim\exp(-cm)\|\mathcal{K}^{\ell}_{T^{0}}v\|_{V(\mathsf{N}^{\ell}(T^{0}))},\quad v\in V^{0},$
@@
We apply Theorem 4.1 with $\mathcal{K}_{T^{0}}$, $V$, and $\mathsf{N}^{\ell}(T^{0})$ replaced by $\mathcal{K}_{T^{0}}^{\ell}$, $V_{T^{0}}^{\ell}$, and $\mathsf{N}^{m}(T^{0})$, respectively. The proof carries over without modification. ∎
@@
(5.9) $\|\mathcal{K}_{T^{0}}^{\ell}v\|^{2}_{V}$ $\lesssim a(\mathcal{K}_{T^{0}}^{\ell}v,\mathcal{K}_{T^{0}}^{\ell}v)=-a_{T^{0}}(\mathcal{I}_{H}v,\mathcal{K}_{T^{0}}^{\ell}v)-b_{T^{0}}(v-\mathcal{I}_{H}v,\lambda^{\ell}).$

### v2

@@ Abstract.
Institute for Applied and Numerical Mathematics, Karlsruhe Institute of Technology, Englerstr. 2, 76131 Karlsruhe, Germany, (moritz.hauck@kit.edu)
Department of Mathematical Sciences, Chalmers University of Technology and University of Gothenburg, Göteborg 412 96, Sweden, (axel@chalmers.se, malinni@chalmers.se)
@@ 1. Introduction
From a numerical perspective, solving mixed-dimensional PDEs presents several challenges. Classical finite element methods require meshes that resolve the complex geometry of embedded features, which can become intractable for a large number of interfaces. The situation is further complicated by the presence of strongly heterogeneous and oscillatory coefficients, which must also be resolved by the considered meshes. This situation is typical in porous media flow where the permeability varies rapidly in the bulk regions between interfaces. This paper focuses precisely on such a setting: we study a mixed-dimensional, elliptic diffusion problem with highly heterogeneous coefficients. For such problems, we develop a multiscale method that constructs problem-adapted, locally computable basis functions defined on a coarse mesh. Since these basis functions capture the essential features of the underlying problem, accurate approximations can be obtained even on very coarse meshes that do not resolve the fine-scale features of the problem.
In general, numerous numerical methods have been developed to solve partial differential equations on surfaces. A foundational contribution to the finite element approximation of the Laplace–Beltrami equation was given in [13], where the method is formulated on a polyhedral approximation of the surface (see also the comprehensive overview in [12]). Alternative approaches include trace-based methods [8, 7, 33], in which surface functions are represented as traces of functions defined in a higher-dimensional ambient space. Finite element methods for coupled bulk–interface problems have also been the subject of extensive research; see, for example, [16, 32]. More recently, a fitted finite element method has been introduced in [25], with particular emphasis on establishing Poincaré-type inequalities for the coupled bulk–interface model.
@@
The remainder of the paper is organized as follows. In section 2, we introduce the mixed-dimensional model problem. A prototypical multiscale method that achieves optimal convergence rates under minimal structural assumptions is presented in section 3. In section 4, we establish the exponential decay of the corresponding prototypical basis functions, motivating their localization, which is discussed in section 5. Based on the resulting localized basis functions, a practical multiscale method is developed in section 6. In section 7, we extend the methodology to allow for interfaces that vary of the fine computational mesh. Finally, in section 8, we present a series of numerical experiments that support our theoretical findings.
@@ 2.1. Geometry
To formulate assumptions on the topology of the problem, we define, for each codimension $c$, a topological subspace $X^{c}\subset\mathbb{R}^{d}$. For codimension $c=0$, we set $X^{0}=\penalty 10000\ \Omega$, and for codimensions $c>0$, we define $X^{c}=X^{c-1}\setminus\Omega^{c-1}.$ Standard topological concepts such as openness, closedness, and density are understood with respect to the subspace topology on $X^{c}$. The closure, boundary, and interior operators are denoted by $\overline{\cdot}$, $\partial\cdot$, and $\operatorname{int}(\cdot)$, respectively. We assume that each subdomain segment $\Omega^{c}_{\ell}$, for $\ell\in L_{c}$, is open in $X^{c}$, and that their union $\Omega^{c}$ is dense in $X^{c}$. Furthermore, we suppose that for any pair $(i,j)\in L_{c}\times L_{c+1}$, we either have the inclusion $\Omega^{c+1}_{j}\subseteq\partial\Omega^{c}_{i}$ or it holds $\Omega^{c+1}_{j}\cap\overline{\Omega^{c}_{i}}=\emptyset$. This condition allows us to define an adjacency relation $E^{c}$ between segments of codimensions $c$ and $c+1$ as
@@ 2.2. Function spaces
using the classical notation for product spaces. Here, the space $V^{0}$, defined on the bulk domain of codimension 0, is broken along the interfaces between subdomain segments. Specifically we let
@@ 2.3. Problem formulation
$-\nabla^{1}\cdot\left(A^{1}\nabla^{1}u^{1}\right)-\sum_{(i,j)\in E^{0}}B^{1}\!\left(u^{0}_{i}-u^{1}_{j}\right)$ $=f^{1},$ $\text{in }\Omega^{1},$
@@
The coercivity proof employs an iterative procedure, starting by bounding norms on subdomain segments far from the domain boundary and progressively extending these bounds to neighboring segments until reaching those where a Friedrichs-type inequality applies; see [25, Lem. 1] for details. The boundedness can be proved using a standard trace inequality; see [25, Lem. 2]. ∎
@@ 3. Prototypical method
In this section, we present a prototypical multiscale method that achieves optimal-order approximations without any pre-asymptotic effects, under minimal structural assumptions on the coefficients. The LOD method is typically defined with respect to a coarse mesh of the domain. In the present mixed-dimensional setting, coarse meshes are required for both the bulk domain and the interface domain, denoted by $\mathcal{T}_{H}^{0}$ and $\mathcal{T}_{H}^{1}$, respectively. They are henceforth referred to as the bulk mesh and the interface mesh, respectively. Note that the considered coarse meshes are not required to resolve the fine-scale variations of the coefficients. In general, there is considerable flexibility in how these meshes are chosen. Even complex-shaped elements representing fractures can be accommodated, provided the elements are non-degenerate and a local element-wise Poincaré inequality holds.
@@
(3.5b) $\qquad\qquad b(\tilde{\varphi}_{T^{c}},\mu)$ $=\quad$ $\mu_{T^{c}},$ $\text{ for all }\mu\in\mathbb{R}^{N},$
@@
(3.7) $\begin{split}\|\nabla^{c}v_{T^{c}}\|_{L^{2}(T^{c})}&\lesssim H^{-1}\|v_{T^{c}}\|_{L^{2}(T^{c})}\lesssim H^{(d-c)/2-1},\quad c\in\{0,1\}.\end{split}$
@@
We begin by establishing well-posedness of problem 3.5 by verifying the inf–sup stability of the bilinear form $b$. Specifically, we show that there exists a constant $C>0$ such that
@@
Given any $\mu\in\mathbb{R}^{N}$, we construct a function $v\in V_{0}$ componentwise as
@@
Having introduced the prototypical basis functions, we can define a projection operator $\mathcal{R}\colon V\to\tilde{V}_{H}$ that preserves the QOI from 3.1 by
@@
This operator coincides with the $a$-orthogonal projection onto $\tilde{V}_{H}$, since for any $v\in\penalty 10000\ V_{0}$ and $w\in\tilde{V}_{H}$ it holds that $a(v-\mathcal{R}v,w)=0$, noting that $v-\mathcal{R}v\in W$. This orthogonality relation also implies the continuity of $\mathcal{R}$, i.e.,
@@ 4. Exponential decay of prototypical basis
where $\mathcal{T}^{c}(z^{c})$ denotes the set of elements in $\mathcal{T}_{H}^{c}$ sharing the node $z^{c}$, and $\Lambda_{z^{c}}$ is the classical hat function associated with node $z^{c}$. For boundary nodes, the corresponding nodal values of $\mathcal{I}_{H}^{c}v$ are set to zero to conform to the prescribed Dirichlet boundary conditions. For $c\in\{0,1\}$, this operator satisfies the following local approximation and stability estimates:
(4.2) $H^{-1}\|v-\mathcal{I}_{H}^{c}v\|_{L^{2}(T^{c})}+\|\nabla^{c}\mathcal{I}_{H}^{c}v\|_{L^{2}(T^{c})}\lesssim\|\nabla^{c}v\|_{L^{2}(\mathsf{N}^{c}(T^{c}))},\;v\in V^{c},\;T^{c}\in\mathcal{T}_{H}^{c}$
@@
where $a_{T^{0}}$ and $b_{T^{0}}$ denote, for any $T^{0}\in\mathcal{T}_{H}^{0}$, suitable restrictions of the bilinear forms $a$ and $b$, satisfying the properties
@@
(4.7) $\begin{split}a_{T^{0}}(v,w)&\coloneqq({A^{0}\nabla^{0}v^{0}},{\nabla^{0}w^{0}})_{T^{0}}+\sum_{T^{1}\subset\partial T^{0}}\frac{1}{n(T^{1})}({A^{1}\nabla^{1}v^{1}},{\nabla^{1}w^{1}})_{T^{1}}\\ &\qquad+\sum_{T^{1}\subset\partial T^{0}}({B^{1}(v^{0}|_{T^{0}}-v^{1})},{w^{0}|_{T^{0}}-w^{1}})_{T^{1}},\end{split}$
@@
(4.9) $\begin{split}\|v\|_{V(S)}^{2}&\coloneqq\|\nabla^{0}v^{0}\|_{L^{2}(S^{0})}^{2}+\|\nabla^{1}v^{1}\|_{L^{2}(S^{1})}^{2}+\hskip-5.38193pt\sum_{\begin{subarray}{c}\{T^{1}\subset\partial T^{0}\,:\,\\ T^{0}\subset S^{0},\,T^{1}\subset S^{1}\}\end{subarray}}\hskip-5.38193pt\|v^{0}|_{T^{0}}-v^{1}\|_{L^{2}(T^{1})}^{2}.\end{split}$
@@
(4.10) $\|\mathcal{K}_{T^{0}}v\|_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}\lesssim\exp(-c\ell)\|\mathcal{K}_{T^{0}}v\|_{V(\Omega)},\quad v\in V_{0}.$
@@
For a given $v\in V_{0}$, define $\varphi\coloneqq\mathcal{K}_{T^{0}}v$. Introduce a cut-off function $\eta\in V_{H}^{0}$, defined on the bulk domain, and uniquely characterized by the following properties:
@@
$|\lambda_{K^{0}}|$ $=|b(\lambda,b_{K^{0}})|=|a(\varphi,b_{K^{0}})|=|(A^{0}\nabla^{0}\varphi^{0},\nabla^{0}v_{K^{0}})_{L^{2}(\Omega^{0})}|$
@@
(4.18) $H^{-d+3}\sum_{K^{1}\subset R_{T^{0}}^{1}}|\lambda_{K^{1}}|^{2}\lesssim\|\nabla^{0}\varphi^{0}\|^{2}_{L^{2}(\mathsf{N}^{0}(R_{T^{0}}^{0}))}+\|\nabla^{1}\varphi^{1}\|^{2}_{L^{2}(R_{T^{0}}^{1})}.$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}\leq C\|\varphi\|^{2}_{V(\mathsf{N}_{\ell+1}(T^{0})\setminus\mathsf{N}_{\ell-2}(T^{0}))}.$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}$ $\leq\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell-1}(T^{0}))}\leq C\|\varphi\|^{2}_{V(\mathsf{N}_{\ell}(T^{0})\setminus\mathsf{N}_{\ell-3}(T^{0}))}$
@@
$\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell}(T^{0}))}\leq\gamma\|\varphi\|^{2}_{V(\Omega\setminus\mathsf{N}_{\ell-3}(T^{0}))}\leq\gamma^{\lfloor\ell/3\rfloor}\|\varphi\|^{2}_{V(\Omega)},$
@@ 5. Localization
$\|\mathcal{K}^{\ell}_{T^{0}}v\|_{V(\Omega\setminus\mathsf{N}_{m}(T^{0}))}\lesssim\exp(-cm)\|\mathcal{K}^{\ell}_{T^{0}}v\|_{V(\mathsf{N}_{\ell}(T^{0}))},\quad v\in V^{0},$
@@
We apply theorem 4.1 with $\mathcal{K}_{T^{0}}$, $V$, and $\mathsf{N}_{\ell}(T^{0})$ replaced by $\mathcal{K}_{T^{0}}^{\ell}$, $V_{T^{0}}^{\ell}$, and $\mathsf{N}_{m}(T^{0})$, respectively. The proof carries over without modification. ∎
@@
(5.9) $\|\mathcal{K}_{T^{0}}^{\ell}v\|^{2}_{V}$ $\lesssim a(\mathcal{K}_{T^{0}}^{\ell}v,\mathcal{K}_{T^{0}}^{\ell}v)=a_{T^{0}}(\mathcal{I}_{H}v,\mathcal{K}_{T^{0}}^{\ell}v)+b_{T^{0}}(v-\mathcal{I}_{H}v,\lambda^{\ell}).$